OpenStudy (anonymous):

Match each item in Column A to an answer in Column B. Are the functions linear or nonlinear? Column A Column B 1. y - 4 = -8(x-1) 2. y = x4 3. y - x2 = 4.5 4. 3x + 5y = 15 A. linear B. nonlinear

OpenStudy (anonymous):

@princeharryyy

OpenStudy (princeharryyy):

1 linear

OpenStudy (princeharryyy):

and 4 is linear rest are non linear

OpenStudy (anonymous):

When is the function constant? from x = to x = A graph of a function. The function has three sections. Section 1 is a line from point 0, 35 to point 2, 15. Section 2 is a line from point 2, 15 to point 4, 15. Section 3 is a line from point 4, 15 to point 5, 5. http://static.k12.com/calms_media/media/1524000_1524500/1524414/1/bacd1c3f315f11bf0f0f3ffcacf0e57b17950ca1/MS_IMC-140901-106147.jpg

OpenStudy (anonymous):

@princeharryyy

OpenStudy (princeharryyy):

0,35 to 5,5

OpenStudy (princeharryyy):

evry where

OpenStudy (anonymous):

Im sorry i didnt understand that

OpenStudy (princeharryyy):

sorry! I didn't read it carefully, cos I was talkin to some1 so, sorry.. ans is 2,15 to 4,15

OpenStudy (anonymous):

Thank you lol

OpenStudy (princeharryyy):

yw

OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the points (2, 5) and (3, 0). y − = (x − 2)

OpenStudy (anonymous):

The ? marks are - signs

OpenStudy (princeharryyy):

y-5 = -5(x-2)

OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the point (−1, 2) and has a slope of 0.9. y = x +

OpenStudy (princeharryyy):

y = 0.9x + 2.9

OpenStudy (anonymous):

A cleaning company charges a fee of $65 in addition to $10.50 per hour of cleaning. Let x represent the number of hours cleaning, and let y represent the total charge. Complete the function equation that relates the total charge to the number of hours spent cleaning. y = x +

OpenStudy (anonymous):

This is last one

OpenStudy (princeharryyy):

y = 10.50x + 65

OpenStudy (anonymous):

Thx I'll tell u my grade i lil bit

OpenStudy (princeharryyy):

ok

OpenStudy (anonymous):

80% !!!!!!!!!!!!!!!!!!!! Thank you!!!!!!!!!

OpenStudy (princeharryyy):

yw

OpenStudy (princeharryyy):

bieee, i need to go :) have a nice day!

OpenStudy (anonymous):

U to!

OpenStudy (princeharryyy):

its night 2 am. well tnku

OpenStudy (princeharryyy):

@*louisa*